Express the following numbers in IEEE 32-bit floating-point format:

a=-8 b=-7 c=-2.5 d=-1/4
Answer #1

0 Given numbers 9:-8; 6:-1; C8-2.5 and d =-74 all axe negative numbers, we need to find JEEE 33-4t floating - peint format t0000 0000 0000 0000 2 Steps: Normalize the mantissa for that a) Semove the leading (the leffurost) (the leffmost) bit, since3 > - Co-2-5 sed 0.5 X2 = lood, Step 28 - Q.5) = (10.1), 10.1 Stepa - (10.12 Sign=1 Step 3 - we have Exponent (un adquitect)